# Bibliography1.bib

### v2

@@ Abstract.
We investigate, under a volume constraint and among sets contained in a Euclidean half-space, the minimization problem of an energy functional given by the sum of a capillarity perimeter, a nonlocal repulsive term and a gravitational potential energy. The capillarity perimeter assigns a constant weight to the portion of the boundary touching the boundary of the half-space. The nonlocal term is represented by a double integral of a positive kernel $g$, while the gravitational term is represented by the integral of a positive potential $G$.
We first establish existence of volume-constrained minimizers in the small mass regime, together with several qualitative properties of minimizers. The existence result holds even for rather general choices of kernels in the nonlocal term, including attractive-repulsive ones.
@@
Finally, we prove a generalized existence result of minimizers holding for all masses, meaning that the infimum of the problem is realized by a finite disjoint union of sets thought located at “infinite distance” one from the other.
@@ 1. Introduction
where $\mathcal{H}^{k}$, with $k\geq 0$, denotes the $k$-dimensional Hausdorff measure in $\mathbb{R}^{n}$, and $\partial^{*}E$ denotes the reduced boundary of $E$, see Section 2. Interpreting the perimeter as a measure of the surface tension of a liquid drop, the constant $\lambda$ basically represent the relative adhesion coefficient between a liquid drop and the solid walls of the container given by $\{x_{n}>0\}$. If $m>0$, minimizers for the isoperimetric capillarity problem
@@
The first result in this paper is an existence result in the capillarity context and in the small mass regime, together with a qualitative properties of volume constrained minimizers.
@@
Let $g$ be a $\mathscr{R}$-admissible $q$-decreasing function, $q\geq 0$, and let $G$ be a $\mathscr{G}$-admissible function. There exists a mass $\bar{m}=\bar{m}(n,\lambda,g,G,q)>0$ such that, for every $m\in(0,\bar{m})$, there exists a minimizer of $\mathscr{F}^{\lambda}$ in the class
@@
Finally, if $g$ is also 00-decreasing, infinitesimal and symmetric and $G$ is coercive, minimizers have no holes, i.e., if $E$ is a minimizer of $\mathscr{F}^{\lambda}$ in $\mathcal{A}_{m}$, there is no set $F\subset\mathbb{R}^{n}\setminus(H\cup E)$ with $|F|>0$ such that
@@
Let us make some comments on the definitions present in Theorem 1.1, while referring to Section 2 for their precise enunciation. The “admissibility” requirements on the kernels just refer to some necessary integrability conditions. The infinitesimality of $g$ and the coercivity of $G$ concern the behavior of these functions as the variable diverge, while the symmetry of $g$ is referred to the symmetry with respect to the origin. The $q$-decreasing property is a way more general condition than the classical radial decreasing condition, denoted in this setting as 00-decreasing property. We point out that attractive-repulsive kernels of the type
@@
are $q$-decreasing for any $q\geq\beta_{1}$, even if they diverge positively as $|x|\to+\infty$, see Definition 2.2 and Remark 2.4. In particular attractive-repulsive kernels as in (1.3) represent a possible choice in the definition of $\mathscr{F}^{\lambda}$ in Theorem 1.1. Minimization problems for attractive-repulsive functionals have been widely studied in the last years. Existence and nonexistence results are addressed in [BurchardChoksiTopaloglu, FrankLiebLiquidSolid, FrankLiebProof], while stability and uniqueness of minimizers have been respectively studied in [BonaciniCristoferiTopaloglu, LopesUniqueness]. We refer to [CarazzatoNote, CarazzatoPratelli, CarazzatoPratelliTopaloglu, CarrilloDelgadinoMellet] for further results about analogous problems.
We remark that, by a symmetry argument, analyzing the Euler-Lagrange equation of problem (1.2), it is possible to verify that the sets $B^{\lambda}(m,x)$ are not volume constrained minimizers of $\mathscr{F}^{\lambda}$; actually, the isoperimetric bubbles $B^{\lambda}(m,x)$ are not even volume constrained critical points of $\mathscr{F}^{\lambda}$. It is left as a future project to study quantitative properties of minimizers to  (1.2), such as the proximity of minimizers from bubbles $B^{\lambda}(m,x)$ in terms of the smallness of the mass.
For large masses and for suitable choices of $g$, the repulsive interaction dominates and the variational problem in Theorem 1.1 does not admit a minimizer.
@@
Therefore, for a general $g$, existence may fail for masses large enough, since minimizers tend to split in two or more components which then move apart one from the other in order to decrease the nonlocal energy. To capture this phenomenon, it is convenient to introduce a generalized energy defined as
@@
Let $g$ be a $\mathscr{R}$-admissible $q$-decreasing function, $q\geq 0$, and let $G$ be a $\mathscr{G}$-admissible function. For every $m>0$ there exists a minimizer of ~$\tilde{\mathscr{F}}^{\lambda}$ in the class
@@
~$\tilde{\mathscr{F}}^{\lambda}(E)=\sum_{j=1}^{h}\mathscr{F}^{\lambda}(E^{j})=\inf\left\{\tilde{\mathscr{F}}^{\lambda}(\Omega)\ :\ \Omega\in\mathcal{A}\right\}.$
@@
Acknowledgments
The author is member of INdAM - GNAMPA. The author is grateful to Marco Pozzetta for many suggestions and for stimulating discussions.
@@ 2.2. Preliminary results on the capillarity functional
$D_{\lambda}(E):=\frac{P_{\lambda}(E)-P_{\lambda}(B^{\lambda}(m))}{P_{\lambda}(B^{\lambda}(m))}.$
Then the following sharp quantitative isoperimetric inequality holds
@@ 2.3. Definitions
We give some definitions for the Riesz-type potential $\mathscr{R}$ and the gravity-type potential $\mathscr{G}$.
@@
A function $g:\mathbb{R}^{n}\setminus\{0\}\to(0,\infty)$ is $\mathscr{R}$-admissible if $\mathscr{R}(B_{1})<\infty$.
A $\mathscr{R}$-admissible function $g:\mathbb{R}^{n}\setminus\{0\}\to(0,\infty)$ is $q$-decreasing, for some $q\in[0,\infty)$, if for every $x\in\mathbb{R}^{n}\setminus\{0\}$ and every $\alpha>1$ it holds
@@ A $\mathscr{R}$-admissible function $g:\mathbb{R}^{n}\setminus\{0\}\to(0,\infty)$ is symmetric if
The functions $\frac{1}{|x|^{\beta}}$, for $\beta\in(0,n)$, are $\mathscr{R}$-admissible, 00-decreasing, infinitesimal and symmetric.
@@
The attractive-repulsive kernels $|x|^{\beta_{1}}+\frac{1}{|x|^{\beta_{2}}}$, for $\beta_{1}>0$ and $\beta_{2}\in(0,n)$, are $\mathscr{R}$-admissible $\beta_{1}$-decreasing symmetric functions. At the same time they diverge positively as $|x|\to+\infty$.
@@
A function $G:(0,\infty)\to(0,\infty)$ is $\mathscr{G}$-admissible if
@@ 3.1. Existence
Let $g$ be a $\mathscr{R}$-admissible $q$-decreasing function and let $G$ be a $\mathscr{G}$-admissible function. There exists a mass $\bar{m}=\bar{m}(n,\lambda,g,G,q)>0$ such that, for all $m\in(0,\bar{m})$, there exists a minimizer of $\mathscr{F}^{\lambda}$ in the class
@@
$\mathscr{R}(\bar{Q}_{1})\geq\sum_{i=0}^{(2N)^{n}}\mathscr{R}(Q_{\frac{1}{2N}}^{i})=(2N)^{n}\mathscr{R}(\bar{Q}_{\frac{1}{2N}}).$
@@
Let $E\subset\mathbb{R}^{n}\setminus H$ be a set of finite perimeter. Let $g$ be $\mathscr{R}$-admissible and $q$-decreasing and let $G$ be $\mathscr{G}$-admissible. If $\alpha>1$, then
@@
Since $g$ is $q$-decreasing, we have
@@
Let $g$ be $\mathscr{R}$-admissible and $q$-decreasing and let $G$ be $\mathscr{G}$-admissible. There exists $\bar{m}>0$, depending on $n$, $\lambda$, $g$, $G$ and $q$, such that, for every $m\in(0,\bar{m})$ and every set of finite perimeter $F\subset\mathbb{R}^{n}\setminus H$ with $|F|=m$, there exists a set of finite perimeter $L$ with
@@
$\begin{split}D_{\lambda}(F)&=\frac{c(n,\lambda)}{m^{\frac{n-1}{n}}}(P_{\lambda}(F)-P_{\lambda}(B^{\lambda}(m)))\\ &\leq\frac{c(n,\lambda)}{m^{\frac{n-1}{n}}}([\mathscr{R}(B^{\lambda}(m))-\mathscr{R}(F)]+[\mathscr{G}(B^{\lambda}(m))-\mathscr{G}(F)])\\ &\leq\frac{c(n,\lambda)}{m^{\frac{n-1}{n}}}(\mathscr{R}(B^{\lambda}(m))+\mathscr{G}(B^{\lambda}(m)))\\ &\leq c(n,\lambda,g,G)\,\left(\frac{m}{|B^{\lambda}|}\right)^{\frac{1}{n}}.\end{split}$
@@
$\alpha_{\lambda}(F)\leq c(n,\lambda)\sqrt{D_{\lambda}(F)}\leq c(n,\lambda,g,G)\left(\frac{m}{|B^{\lambda}|}\right)^{\frac{1}{2n}}$
@@
By Lemma 3.5 there exists a minimizing sequence with uniformly bounded sets. The lower semicontinuity of $P_{\lambda}$ [PascalePozzettaQuantitative, Lemma 3.9] and the continuity of $\mathscr{R}$ and $\mathscr{G}$ under strong $L^{1}$ convergence allow to conclude the proof. ∎

### v3

# Bibliography2.bib

@@ Abstract.
We investigate, under a volume constraint and among sets contained in a Euclidean half-space, the minimization problem of an energy functional given by the sum of a capillarity perimeter, a nonlocal interaction term and a gravitational potential energy. The capillarity perimeter assigns a constant weight to the portion of the boundary touching the boundary of the half-space. The nonlocal term is represented by a double integral of a positive kernel $g$, while the gravitational term is represented by the integral of a positive potential $G$.
We first establish existence of volume-constrained minimizers in the small mass regime, together with several qualitative properties of minimizers. The existence result holds for rather general choices of kernels in the nonlocal interaction term, including attractive-repulsive ones.
@@
Finally, we prove a generalized existence result of minimizers holding for all masses and general nonlocal interaction terms, meaning that the infimum of the problem is realized by a finite disjoint union of sets thought located at “infinite distance” one from the other.
@@ 1. Introduction
where $\mathcal{H}^{k}$, with $k\geq 0$, denotes the $k$-dimensional Hausdorff measure in $\mathbb{R}^{n}$, $H:=\{x_{n}\leq 0\}$ and $\partial^{*}E$ denotes the reduced boundary of $E$, see Section 2. Interpreting the perimeter as a measure of the surface tension of a liquid drop, the constant $\lambda$ basically represent the relative adhesion coefficient between a liquid drop and the solid walls of the container given by $\{x_{n}>0\}$. If $m>0$, minimizers for the isoperimetric capillarity problem
@@
The first result in this paper is an existence result in the capillarity context and in the small mass regime, together with a bound on the Fraenkel asymmetry
$\alpha_{\lambda}(E):=\inf\left\{\frac{|E\Delta B^{\lambda}(m,x)|}{m}\ :\ x\in\{x_{n}=0\}\right\},\qquad{\rm for}\,{\rm any}\;E\subset\{x_{n}>0\}\;{\rm with}\;|E|=m,$
and some qualitative properties of volume constrained minimizers. At the same time, under suitable conditions on the potential energies, existence result extends to all masses.
@@
Let $g$ be a $\mathscr{R}$-admissible $q$-growing function, $q\geq 0$, and let $G$ be a $\mathscr{G}$-admissible function. There exists a mass $\bar{m}=\bar{m}(n,\lambda,g,G,q)>0$ such that, for every $m\in(0,\bar{m})$, there exists a minimizer $E$ of $\mathscr{F}^{\lambda}$ in the class
$\mathcal{A}_{m}:=\{\Omega\subset\mathbb{R}^{n}\setminus H\,{\rm measurable}\ :\ |\Omega|=m\}$
and it satisfies
$\alpha_{\lambda}(E)\leq c(n,\lambda,g,G)\,m^{\frac{1}{2n}}.$
@@
Furthermore, if $g$ is also 00-growing, infinitesimal and symmetric and $G$ is coercive, minimizers have no holes, i.e., if $E$ is a minimizer of $\mathscr{F}^{\lambda}$ in $\mathcal{A}_{m}$, there is no set $F\subset\mathbb{R}^{n}\setminus(H\cup E)$ with $|F|>0$ such that
@@
Finally, if $g$ is $\mathscr{R}$-admissible and coercive and $G$ is $\mathscr{G}$-admissible and coercive, there exists a minimizer of $\mathscr{F}^{\lambda}$ in $\mathcal{A}_{m}$ for any $m>0$.
Let us make some comments on the definitions present in Theorem 1.1, while referring to Section 2 for their precise enunciation. The “admissibility” requirements on the kernels just refer to some necessary integrability conditions. The infinitesimality of $g$ and the coercivity of $g$ and $G$ concern the behavior of these functions as the variable diverge, while the symmetry of $g$ is referred to the symmetry with respect to the origin. The $q$-growing property is satisfied by rather general nonlocal interaction terms, not only by repulsive ones. Indeed, we point out that classical radial decreasing kernels are 00-growing, but at the same time attractive-repulsive kernels of the type
@@
are $q$-growing for any $q\geq\beta_{1}$, even if they diverge positively as $|x|\to+\infty$, see Definition 2.2 and Remark 2.4. In particular attractive-repulsive kernels as in (1.3) represent a possible choice in the definition of $\mathscr{F}^{\lambda}$ in Theorem 1.1. Minimization problems for attractive-repulsive functionals have been widely studied in the last years. Existence and nonexistence results are addressed in [BurchardChoksiTopaloglu, FrankLiebLiquidSolid, FrankLiebProof], while stability and uniqueness of minimizers have been respectively studied in [BonaciniCristoferiTopaloglu, LopesUniqueness]. We refer to [CarazzatoNote, CarazzatoPratelli, CarazzatoPratelliTopaloglu, CarrilloDelgadinoMellet] for further results about analogous problems.
We remark that, by a symmetry argument, analyzing the Euler-Lagrange equation of problem (1.2), it is possible to verify that the sets $B^{\lambda}(m,x)$ are not volume constrained minimizers of $\mathscr{F}^{\lambda}$; actually, the isoperimetric bubbles $B^{\lambda}(m,x)$ are not even volume constrained critical points of $\mathscr{F}^{\lambda}$. It is left as a future project to study quantitative properties of minimizers to  (1.2), such as the proximity of minimizers from bubbles $B^{\lambda}(m,x)$ in terms of the smallness of the mass. Related results, concerning liquid drops or crystals lying in equilibrium under the action of a potential energy, are addressed in [FigalliMaggiShape, MaggiMihaila].
For large masses and for suitable choices of repulsive kernels $g$, the repulsive interaction dominates and the variational problem in Theorem 1.1 does not admit a minimizer.
@@
Therefore, for a general repulsive kernel $g$, existence may fail for masses large enough, since minimizers tend to split in two or more components which then move apart one from the other in order to decrease the nonlocal energy. To capture this phenomenon, it is convenient to introduce a generalized energy defined as
@@
Let $g$ be a $\mathscr{R}$-admissible $q$-growing function, $q\geq 0$, and let $G$ be a $\mathscr{G}$-admissible function. For every $m>0$ there exists a minimizer of ~$\tilde{\mathscr{F}}^{\lambda}$ in the class
@@
~$\tilde{\mathscr{F}}^{\lambda}(E)=\sum_{j=1}^{h}\mathscr{F}^{\lambda}(E^{j})=\inf\left\{\tilde{\mathscr{F}}^{\lambda}(\Omega)\ :\ \Omega\in\mathcal{A}_{m}\right\}.$
@@
Remark 1.4.
We note that, if $g$ is infinitesimal, then
$\inf\left\{\tilde{\mathscr{F}}^{\lambda}(\Omega)\ :\ \Omega\in\mathcal{A}_{m}\right\}=\inf\left\{\mathscr{F}^{\lambda}(\Omega)\ :\ \Omega\in\mathcal{A}_{m}\right\}.$
The proof can be easily adapted by [NovagaPratelli, Lemma 3.4], with attention given to translating the components of a partition without changing the $n$-th component. In this case, every minimizer of $\mathscr{F}^{\lambda}$ is also a minimizer of ~$\tilde{\mathscr{F}}^{\lambda}$. It is left as a future project to study connections between minimizers and generalized minimizers when $g$ is not infinitesimal.
@@ 2.2. Preliminary results on the capillarity functional
$D_{\lambda}(E):=\frac{P_{\lambda}(E)-P_{\lambda}(B^{\lambda}(m))}{P_{\lambda}(B^{\lambda}(m))},$
for any $E\subset\{x_{n}>0\}$ with volume $|E|=m$. Then the following sharp quantitative isoperimetric inequality holds
@@ 2.3. Definitions
We provide some definitions for the kernels of Riesz-type potential $\mathscr{R}$ and gravity-type potential $\mathscr{G}$. In particular, in the following Definition 2.2 and Definition 2.5 we will impose pointwise requirements on the functions $g$, $G$, i.e., it is understood that we fix pointwise defined representatives for the functions $g$, $G$.
@@
A function $g:\mathbb{R}^{n}\setminus\{0\}\to(0,\infty)$ is $\mathscr{R}$-admissible if $g\in L^{1}_{loc}(\mathbb{R}^{n}\setminus\{0\})$ and $\mathscr{R}(B_{1})<\infty$.
A $\mathscr{R}$-admissible function $g:\mathbb{R}^{n}\setminus\{0\}\to(0,\infty)$ is $q$-growing, for some $q\in[0,\infty)$, if for every $x\in\mathbb{R}^{n}\setminus\{0\}$ and every $\alpha>1$ it holds
@@ A $\mathscr{R}$-admissible function $g:\mathbb{R}^{n}\setminus\{0\}\to(0,\infty)$ is symmetric if
A $\mathscr{R}$-admissible function $g:\mathbb{R}^{n}\setminus\{0\}\to(0,\infty)$ is coercive if
$g(x)\to+\infty\qquad{\rm as}\;|x|\to+\infty.$
@@ A $\mathscr{R}$-admissible function $g:\mathbb{R}^{n}\setminus\{0\}\to(0,\infty)$ is coercive if
The functions $\frac{1}{|x|^{\beta}}$, for $\beta\in(0,n)$, are $\mathscr{R}$-admissible, 00-growing, infinitesimal and symmetric.
@@
The attractive-repulsive kernels $|x|^{\beta_{1}}+\frac{1}{|x|^{\beta_{2}}}$, for $\beta_{1}>0$ and $\beta_{2}\in(0,n)$, are $\mathscr{R}$-admissible $\beta_{1}$-growing symmetric functions. At the same time they diverge positively as $|x|\to+\infty$.
@@
A function $G:(0,\infty)\to(0,\infty)$ is $\mathscr{G}$-admissible if $G\in L^{1}_{loc}(0,\infty)$,
@@ 3.1. Existence
Let $g$ be a $\mathscr{R}$-admissible $q$-growing function and let $G$ be a $\mathscr{G}$-admissible function. There exists a mass $\bar{m}=\bar{m}(n,\lambda,g,G,q)>0$ such that, for all $m\in(0,\bar{m})$, there exists a minimizer of $\mathscr{F}^{\lambda}$ in the class
@@
$\mathscr{R}(\bar{Q}_{1})\geq\sum_{i=1}^{(2N)^{n}}\mathscr{R}(Q_{\frac{1}{2N}}^{i})=(2N)^{n}\mathscr{R}(\bar{Q}_{\frac{1}{2N}}).$
@@
Let $E\subset\mathbb{R}^{n}\setminus H$ be a set of finite perimeter. Let $g$ be $\mathscr{R}$-admissible and $q$-growing and let $G$ be $\mathscr{G}$-admissible. If $\alpha>1$, then
@@
Since $g$ is $q$-growing, we have
@@
Let $g$ be $\mathscr{R}$-admissible and $q$-growing and let $G$ be $\mathscr{G}$-admissible. There exists $\bar{m}>0$, depending on $n$, $\lambda$, $g$, $G$ and $q$, such that, for every $m\in(0,\bar{m})$ and every set of finite perimeter $F\subset\mathbb{R}^{n}\setminus H$ with $|F|=m$, there exists a set of finite perimeter $L$ with
@@
(3.3) $\begin{split}D_{\lambda}(F)&=\frac{c(n,\lambda)}{m^{\frac{n-1}{n}}}(P_{\lambda}(F)-P_{\lambda}(B^{\lambda}(m)))\\ &\leq\frac{c(n,\lambda)}{m^{\frac{n-1}{n}}}([\mathscr{R}(B^{\lambda}(m))-\mathscr{R}(F)]+[\mathscr{G}(B^{\lambda}(m))-\mathscr{G}(F)])\\ &\leq\frac{c(n,\lambda)}{m^{\frac{n-1}{n}}}(\mathscr{R}(B^{\lambda}(m))+\mathscr{G}(B^{\lambda}(m)))\\ &\leq c(n,\lambda,g,G)\,\left(\frac{m}{|B^{\lambda}|}\right)^{\frac{1}{n}}.\end{split}$
@@
(3.4) $\alpha_{\lambda}(F)\leq c(n,\lambda)\sqrt{D_{\lambda}(F)}\leq c(n,\lambda,g,G)\left(\frac{m}{|B^{\lambda}|}\right)^{\frac{1}{2n}}$
@@
Remark 3.6.
Arguing exactly as in (3.2), (3.3) and (3.4), we easily deduce that if $E$ is a minimizer of $\mathscr{F}^{\lambda}$ in $\mathcal{A}_{m}$ with $m$ sufficiently small then
$\alpha_{\lambda}(E)\leq c(n,\lambda,g,G)m^{\frac{1}{2n}}.$
@@
By Lemma 3.5 there exists a minimizing sequence with uniformly bounded sets. The lower semicontinuity of $P_{\lambda}$ [PascalePozzettaQuantitative, Lemma 3.7] and the continuity of $\mathscr{R}$ and $\mathscr{G}$ under strong $L^{1}$ convergence (which holds by dominated convergence theorem and uniformly boundedness of the minimizing sequence) allow to conclude the proof. ∎
The following proposition states that, if the nonlocal kernel $g$ and the gravitational term $G$ are coercive, we have existence of minimizers for all values of $m$.
Proposition 3.7.
Let $g$ be a $\mathscr{R}$-admissible coercive function and let $G$ be a $\mathscr{G}$-admissible coercive function. Then, for every $m>0$, there exists a minimizer of $\mathscr{F}^{\lambda}$ in the class
Proof.
Let us consider a minimizing sequence $\left\{E_{i}\right\}_{i\in\mathbb{N}}$ of $\mathscr{F}^{\lambda}$ in $\mathcal{A}_{m}$. In particular
$\sup_{i\in\mathbb{N}}\mathscr{F}^{\lambda}(E_{i})=C<+\infty.$
By [PascalePozzettaQuantitative, Corollary 2.5] it holds
$P(E_{i})\leq\frac{2}{1-\lambda}P_{\lambda}(E_{i})\leq\frac{2}{1-\lambda}\mathscr{F}^{\lambda}(E_{i})\leq C,\qquad\forall i\in\mathbb{N}.$
By [AntonelliNardulliPozzetta, Lemma 2.10] and [LeonardiRitoreVernadakis, Corollary 3.25] we get the existence of a constant $\bar{c}=\bar{c}\left(m,n,\sup_{i}P(E_{i})\right)$ $>0$ such that for every $i\in\mathbb{N}$ there exists $x_{i}\in\left\{x_{n}>0\right\}$ with
$\left|E_{i}\cap B_{1}(x_{i})\right|\geq\bar{c}.$
By the coercivity of $G$ we have that $\left(x_{i}\right)_{n}$ is uniformly bounded. Indeed, if by contradiction $\left(x_{i}\right)_{n}\to+\infty$, then
$C\geq\int_{E_{i}}G(x_{n})\,{\rm d}x\geq\int_{E_{i}\cap B_{1}(x_{i})}G(x_{n})\,{\rm d}x\geq\bar{c}\left(\inf_{B_{1}(x_{i})}G(x_{n})\right)\to+\infty.$
By the lower semicontinuity of $\mathscr{F}^{\lambda}$ under strong $L^{1}$ convergence, the existence of a minimizer follows if we show that for every $\varepsilon>0$ and $i\in\mathbb{N}$ there exists $R>0$ with
(3.12) $\sup_{i\in\mathbb{N}}\left|E_{i}\setminus B_{R}(x_{i})\right|<\varepsilon.$
In order to prove (3.12), note that
$\begin{split}C\geq\mathscr{R}(E_{i})&\geq\int_{E_{i}}\int_{E_{i}\setminus B_{R}(x)}g(y-x)\,{\rm d}y\,{\rm d}x\geq\left(\inf_{|x|>R}g(x)\right)\int_{E_{i}}\left|E_{i}\setminus B_{R}(x)\right|\,{\rm d}x\\ &\geq\left(\inf_{|x|>R}g(x)\right)\int_{E_{i}\cap B_{1}(x_{i})}\left|E_{i}\setminus B_{R}(x)\right|\,{\rm d}x\geq\bar{c}\left(\inf_{|x|>R}g(x)\right)\left|E_{i}\setminus B_{R+1}(x_{i})\right|\end{split}$
and
$\left|E_{i}\setminus B_{R+1}(x_{i})\right|\leq\frac{C}{\bar{c}}\left(\inf_{|x|>R}g(x)\right)^{-1}\to 0\qquad{\rm as}\quad R\to+\infty.$